OpenStudy (anonymous):

Please help me check my work. I needed the equation for the tangent line to the elipse 2x^2 + y^2 = 3 and that was parallel to 2x + y = 5.

OpenStudy (anonymous):

I first wrote it like this: r: (x,y) = (x0,y0) + lambda v, where v gradient f(x0,y0) = 0. gradient of f(x,y) = (4x,2y) gradient of f(x0,y0) = (4x0,2y0)

OpenStudy (anonymous):

Then I did the same with 2x + y = 5 r2: (x,y) = (x1,y1) + t u gradient of f(x,y) = (2,1) gradient of f(x1,y1) = (2,1) u = (-1,2) v is parallel to (2,1), so for some value of h, v = h (-1,2) v = (2y0,-4x0) 2y0 = - h -4x0 = 2h y0 = -h/2 x0 = -h/2 h = -2 y0 = 1 x0 = 1

OpenStudy (anonymous):

r: (x,y) = (1,1) + lambda (2,-4)

OpenStudy (anonymous):

I'm not sure if it's right though...